OpenStudy (anonymous):

Find the surface area of spheres having the following measurements. a.r=6in b.d=24m c.C=7pie cm

OpenStudy (anonymous):

4(pi)r^2

OpenStudy (anonymous):

Is that the formula which I have. I just don't get how to solve it.

OpenStudy (anonymous):

a: \[4\pi(6^{2})\] b: radius is half of diameter. radius is 12. \[4\pi(12^2)\] c: \[C =2\pi r\] \[7\pi=2\pi r\] \[r = \frac{7\pi}{2\pi}\] \[r = \frac{7}{2}\pi\] the surface area is....... \[4\pi(\frac{7}{2}\pi)^2\]

OpenStudy (anonymous):

okay sorry. The surface area for c. is actually \[4\pi(\frac{7}{2})^2\] mistake.

OpenStudy (anonymous):

Thank You Ian!